# DriveMoE: Mixture-of-Experts for Vision-Language-Action Model in End-to-End Autonomous Driving

### v1

@@ 3.4 Skill-Specialized Action MoE
Figure 4: The Skill-Specialized Action Mixture-of-Experts.
Human drivers fluidly transition among different driving skills—such as smoothly cruising down a highway, carefully merging into traffic, swiftly overtaking slower vehicles, or urgently braking in response to sudden obstacles. Each of these driving skills is associated with distinct behavioral patterns and trajectory characteristics. Although the original flow-matching decoder of $\pi_{0}$ could already generate diverse trajectories, employing one single model inevitably averages across these diverse behaviors black2024pi_0 , making the model fail to accurately generate rare yet safety-critical maneuvers.
To address these issues, inspired by human intuition—where drivers naturally select the appropriate driving skill based on the current context, we propose a Skill-Specialized Action MoE architecture built on the original flow-matching trajectory transformer. The central idea is to decompose the policy’s representation of behaviors by replacing each dense feed-forward network (FFN) in the decoder with a Mixture-of-Experts (MoE) layer containing multiple skill-specific experts. Essentially, each decoder layer is no longer a single monolithic mapping, but an ensemble of $K$ expert FFNs each intended to specialize in a subset of driving skills. By conditionally routing each input through a small subset of these experts, the model isolates distinct behavior modes instead of forcing them into a single decoder stream. This design prevents the averaging effect observed in one single model and thereby allocates dedicated model capacity to rare maneuvers. The result is a policy network that preserves the multimodality of the trajectory data, modeling both frequent and infrequent behaviors with appropriate specialization.
@@
The updated feature combines the outputs of individual experts weighted by the router’s confidence:
@@
In practice, we use a sparse activation mechanism jiang2024mixtral to select only a few experts with the highest ranking for calculation (only activate the Top-1 or Top-2 experts), thereby reducing the amount of calculation, preventing mutual interference between experts, and strengthening the degree of expert skill specialization. This sparse routing mechanism is consistent with the mechanism we use in the Vision MoE module, ensuring that each expert clearly focuses on a specific behavior mode.
To explicitly guide our model toward meaningful skill specialization—mirroring structured and intuitive human-defined skill categories—we utilize driving skill labels $y_{k}\in\{1,\dots,K\}$, annotated based on scenarios, and train the skill router via a cross-entropy loss as well:
@@
Additionally, we optimize the entire Action MoE module using a flow-matching trajectory loss $\mathcal{L}_{\text{FM}}$ to ensure accurate trajectory predictions, and introduce a load-balancing regularization loss $\mathcal{L}_{\text{LB}}$ to maintain balanced expert utilization, preventing expert collapse:
@@
where $\lambda_{1}$ represents loss weight of flow matching policy, $\lambda_{2}$ represents loss weight of action router. We introduce noise into the action router following dai2024deepseekmoe , which increases randomness and encourages exploration, effectively mitigating the risk of expert collapse.
@@ 3.5 Two Stage Training: From Teacher-Forcing to Adaptive Training
DriveMoE loads the Paligemma VLM pretrained weights beyer2024paligemma and we finetune it in the autonomous driving scene via a two-stage training procedure. In the first stage, both vision and action MoEs only select ground-truth experts while the router is trained jointly, which significantly stabilize the training. In the second stage, we transition to select experts based on the outputs of Vision and Action MoE routers, removing reliance on GT annotation about experts. It develops robustness against potential mistakes or inaccuracies made by routers, thereby enhancing the overall model’s generalization capability under realistic inference conditions.
@@ 4.1 Datasets & Benchmark & Metric
We employ the CARLA simulator dosovitskiy2017carla (version 0.9.15.1) for closed-loop driving performance evaluation, and adopt the latest public closed-loop evaluation benchmark, Bench2Drive jia2024bench which includes 220 short routes with one challenging corner case per route for analysis of different driving abilities. It provides an official training set, where we use the base set (1000 clips, 950 training, 50 test/validation) for fair comparison with all the other baselines.
We use the official 220 routes and official metrics of Bench2Drive for evaluation. The Driving Score (DS) is defined as the product of Route Completion and Infraction Score, capturing both task completion and rule adherence. The Success Rate (SR) measures the percentage of routes completed successfully within the allocated time and without committing any traffic violations. Efficiency quantifies the vehicle’s velocity relative to surrounding traffic, encouraging progressiveness without aggression. Comfort reflects the smoothness of the driving trajectory. Meanwhile, Bench2Drive evaluates driving capabilities across multiple critical dimensions, including tasks such as Merging, Overtaking, Emergency Braking, Yielding, and Traffic Signs.
4.2 Implementation Details
Vision Routing Annotations: We introduce additional camera-view importance annotations into the Bench2Drive jia2024bench dataset. This annotation approach is both inexpensive and straightforward, yet it significantly improves model performance through efficient and effective utilization of multi-camera inputs. The details about camera annotation rules refer to Appendix A.
Action Routing Annotations: We maintain skill definitions consistent with Bench2Drive jia2024bench setup. There are five driving skills: Merging, Overtaking, Emergency Brake, Give Way, and Traffic Sign.
@@ 4.2 Implementation Details
Training Stage 1. We train the model for 10 epochs. The Vision-Language Model (VLM) component is initialized from the pretrained weights of Paligemma-3b-pt-224 beyer2024paligemma . The VLA and Action MoE experts are optimized separately using two optimizers, both configured as follows: learning rate = $5\times 10^{-5}$, and warmup steps enabled. Gradient clipping is applied with a maximum gradient norm of 1.0. Gradient accumulation is used to simulate a batch size of 1024. To balance different loss components effectively, we set the vision router loss weight $\lambda_{0}$ to 0.05, action router loss weight $\lambda_{2}$ to 0.03, flow matching loss weight $\lambda_{1}$ to 1.
Training Stage 2. We continue training for an additional 5 epochs, initializing from the checkpoint obtained at the end of Stage 1. In this stage, input camera views and action experts are dynamically selected based on outputs from the routers. We set the action router loss weight $\lambda_{2}$ to 0.025, emphasizing trajectory learning. Other hyperparameters remain consistent with Stage 1.
4.3 Comparison with State-of-the-Art Works
As shown in Table 2, our proposed method achieves state-of-the-art (SOTA) performance in terms of both driving score and success rate on the Bench2Drive closed-loop benchmark. Specifically, compared to the baseline Drive-$\pi_{0}$, our method improves the driving score by 22.8% and the success rate by 62.1%. On the open-loop metric, our method achieves the lowest L2 error. We observe that diffusion policy-based trajectory prediction significantly reduces L2 errors compared to traditional methods. However, as highlighted in prior studies such as AD-MLP zhai2023ADMLP , TransFuser++ Jaeger2023ICCV , and Bench2Drive jia2024bench , open-loop metrics mainly serve as indicators of model convergence, whereas closed-loop metrics provide a more reliable assessment of true driving performance. Moreover, in the multi-dimensional capability evaluation, as shown in Table 1, our method obtains state-of-the-art results across five key capabilities and their overall average.
Table 1: Performance on Bench2Drive Multi-Ability Benchmark. * denotes expert feature distillation.
Method Venue Ability (%) $\uparrow$
Merging Overtaking Emergency Brake Give Way Traffic Sign Mean
TCP-traj* wu2022trajectoryguided NeurIPS 2022 8.89 24.29 51.67 40.00 46.28 34.22
AD-MLP zhai2023rethinking Arxiv 2023 0.00 0.00 0.00 0.00 4.35 0.87
UniAD-Base hu2023planning CVPR 2023 14.10 17.78 21.67 10.00 14.21 15.55
ThinkTwice* jia2023thinktwice CVPR 2023 27.38 18.42 35.82 50.00 54.23 37.17
VAD jiang2023vad ICCV 2023 8.11 24.44 18.64 20.00 19.15 18.07
DriveAdapter* jia2023driveadapter ICCV 2023 28.82 26.38 48.76 50.00 56.43 42.08
DriveTrans jia2025drivetransformer ICLR 2025 17.57 35.00 48.36 40.00 52.10 38.60
DiffAD wang2025diffad Arxiv 2025 30.00 35.55 46.66 40.00 46.32 38.79
@@ 4.3 Comparison with State-of-the-Art Works
DriveMoE (Ours) - 34.67 40.00 65.45 40.00 59.44 47.91
@@ 4.4 Ablation Study
[…]
Table 2: Results on the Bench2Drive Benchmark. The result includes both Closed-Loop and Open-Loop metrics. * denotes expert feature distillation.
Method Venue Closed-loop Metric Open-loop Metric
DS $\uparrow$ SR(%) $\uparrow$ Efficiency $\uparrow$ Comfort $\uparrow$ Avg. L2$\downarrow$
TCP-traj* wu2022trajectoryguided NeurIPS 2022 59.90 30.00 76.54 18.08 1.70
AD-MLP zhai2023rethinking Arxiv 2023 18.05 0.00 48.45 22.63 3.64
VAD jiang2023vad ICCV 2023 42.35 15.00 157.94 46.01 0.91
UniAD-Base hu2023planning CVPR 2023 45.81 16.36 129.21 43.58 0.73
ThinkTwice* jia2023thinktwice CVPR 2023 62.44 31.23 69.33 16.22 0.95
DriveAdapter* jia2023driveadapter ICCV 2023 64.22 33.08 70.22 16.01 1.01
GenAD zheng2024genad ECCV 2024 44.81 15.90 - - -
DriveTrans jia2025drivetransformer ICLR 2025 63.46 35.01 100.64 20.78 0.62
MomAD song2025don CVPR 2025 44.54 16.71 170.21 48.63 0.82
WoTE li2025end Arxiv 2025 61.71 31.36 - - -
DiffAD wang2025diffad Arxiv 2025 67.92 38.64 - - 1.55
[…]
DriveMoE(Ours) - 74.22 48.64 175.96 15.31 0.38
Table 3: Drive-$\pi_{0}$ vs DriveMoE. Evaluate the Vision MoE and Action MoE. ”w/o” denotes removing the respective modules from DriveMoE.
[…]
Drive-$\pi_{0}$ 60.45 30.00
w/o Vision MoE 68.68 42.45
w/o Action MoE 67.31 40.56
DriveMoE 74.22 48.64
[…]
5 73.81 47.73
6 74.22 48.64
13 70.88 44.50
44 68.22 43.18
@@
Vision MoE. As shown in Table 5, we investigate the contribution of camera view selection and supervision signals within our Vision MoE module. The baseline (①, Drive-$\pi_{0}$) utilizes two consecutive front-view images ($\boldsymbol{I}_{t}^{\text{front}}+\boldsymbol{I}_{t-1}^{\text{front}}$) primarily to estimate velocities of surrounding agents. Adding a third fixed view such as the back view (②), front-left view (③), or front-right view (④) provides additional spatial context, yielding moderate improvements. By introducing dynamically selected views without supervision (⑤), the driving score and success rate significantly improve. Ultimately, incorporating explicit supervision signals (⑥, DriveMoE) for the dynamic view selection further enhances both driving score and success rate, demonstrating the effectiveness of our Vision MoE module in leveraging dynamic and supervised multi-view perception. Table 6 shows the accuracy of the vision and action routers on the test set under open-loop evaluation.
Table 5: Ablation Study of Vision MoE. Compare different camera view combinations and supervision signals. Dynamic View represents the camera view dynamically selected by the vision router as the top-1 relevant view. ① is our baseline Drive-$\pi_{0}$ using two consecutive front-view images to model velocities of other traffic agents. ⑥ is DriveMoE, adding a dynamically selected camera view supervised by an explicit supervision signal to enhance perception learning.
[…]
$\boldsymbol{I}_{t}^{\text{front}}$+$\boldsymbol{I}_{t-1}^{\text{front}}$ 60.45 30.00
$\boldsymbol{I}_{t}^{\text{front}}$+$\boldsymbol{I}_{t-1}^{\text{front}}$+$\boldsymbol{I}_{t}^{\text{back}}$ 64.52 32.73
$\boldsymbol{I}_{t}^{\text{front}}$+$\boldsymbol{I}_{t-1}^{\text{front}}$+$\boldsymbol{I}_{t}^{\text{front\_left}}$ 65.38 33.64
$\boldsymbol{I}_{t}^{\text{front}}$+$\boldsymbol{I}_{t-1}^{\text{front}}$+$\boldsymbol{I}_{t}^{\text{front\_right}}$ 63.26 31.82
$\boldsymbol{I}_{t}^{\text{front}}$+$\boldsymbol{I}_{t-1}^{\text{front}}$+ Dynamic View $\times$ 69.71 44.09
$\boldsymbol{I}_{t}^{\text{front}}$+$\boldsymbol{I}_{t-1}^{\text{front}}$+ Dynamic View $\checkmark$ 74.22 48.64
@@
Action MoE. We investigate the impact of the number of non-shared experts within our Action MoE, as shown in Table 4. Specifically, configuration ① corresponds to the original five skills defined by Bench2Drive jia2024bench , while ② introduces an additional expert for the classic ParkingExits scenario, resulting in improved performance. To further analyze the effect of expert specialization, we conducted additional experiments: ③ adds experts targeting several challenging scenarios identified from configuration ②, and ④ assigns a distinct expert to each of the 44 scenarios in Bench2Drive. We observe that excessively increasing the number of experts (③, ④) negatively affects performance due to the induced load imbalance among experts. Thus, an appropriate balance in the number of specialized experts is crucial for optimal driving performance.

### v2

@@ 2.4 Skill-Specialized Action MoE
Human drivers fluidly transition among different driving skills—such as smoothly cruising down a highway, carefully merging into traffic, swiftly overtaking slower vehicles, or urgently braking in response to sudden obstacles. Each of these driving skills is associated with distinct behavioral patterns and trajectory characteristics. Although the original flow-matching decoder of $\boldsymbol{\pi}_{\boldsymbol{0}}$ could already generate diverse trajectories, employing one single model inevitably averages across these diverse behaviors [3], making the model fail to accurately generate rare yet safety-critical maneuvers.
To address these issues, inspired by human intuition—where drivers naturally select the appropriate driving skill based on the current context, we propose a Skill-Specialized Action MoE architecture built on the original flow-matching trajectory transformer. The details about flow matching trajectory loss refer to Appendix 6. The central idea is to decompose the policy’s representation of behaviors by replacing each dense feed-forward network (FFN) in the decoder with a Mixture-of-Experts (MoE) layer containing multiple skill-specific experts. Essentially, each decoder layer is no longer a single monolithic mapping, but an ensemble of $K$ expert FFNs each intended to specialize in a subset of driving skills. By conditionally routing each input through a small subset of these experts, the model isolates distinct behavior modes instead of forcing them into a single decoder stream. This design prevents the averaging effect observed in one single model and thereby allocates dedicated model capacity to rare maneuvers. The result is a policy network that preserves the multimodality of the trajectory data, modeling both frequent and infrequent behaviors with appropriate specialization. As illustrated in Figure 4 and Figure 5, we explore two styles of MoE design: token-level (similar to DeepSeek-MoE) and trajectory-level.
2.4.1 Token-level ActionMoE
In Token-level ActionMoE, each token (corresponding to a specific timestep in the trajectory) independently selects its expert(s) based on its local hidden representation. This strategy allows different experts to specialize in modeling short-horizon temporal dependencies—e.g., acceleration, braking, or turning subtasks—within a single trajectory.
Figure 5: Trajectory-Level Skill-Specialized Action Mixture-of-Experts.
@@ 2.4.1 Token-level ActionMoE
It is worth noting that $\mathbf{h}$ is calculated in the dimension of token num. The updated feature combines the outputs of individual experts weighted by the router’s confidence:
@@
In practice, we use a sparse activation mechanism [30] to select only a few experts with the highest ranking for calculation (only activate the Top-1 or Top-2 experts), thereby reducing the amount of calculation, preventing mutual interference between experts, and strengthening the degree of expert skill specialization. This sparse routing mechanism is consistent with the mechanism we use in the Vision MoE module, ensuring that each expert clearly focuses on a specific behavior mode.
Table 1: Performance on Bench2Drive Multi-Ability Benchmark. *: expert feature distillation.
Method Venue Ability (%) $\uparrow$
Merging Overtaking Emergency Brake Give Way Traffic Sign Mean
TCP-traj* [54] NeurIPS 2022 8.89 24.29 51.67 40.00 46.28 34.22
AD-MLP [61] Arxiv 2023 0.00 0.00 0.00 0.00 4.35 0.87
UniAD-Base [16] CVPR 2023 14.10 17.78 21.67 10.00 14.21 15.55
ThinkTwice* [25] CVPR 2023 27.38 18.42 35.82 50.00 54.23 37.17
VAD [31] ICCV 2023 8.11 24.44 18.64 20.00 19.15 18.07
DriveAdapter* [22] ICCV 2023 28.82 26.38 48.76 50.00 56.43 42.08
DriveTrans [28] ICLR 2025 17.57 35.00 48.36 40.00 52.10 38.60
DiffAD [52] Arxiv 2025 30.00 35.55 46.66 40.00 46.32 38.79
[…]
2.4.2 Trajectory-level ActionMoE
In Trajectory-level ActionMoE, instead of routing per token, the router assigns experts at the trajectory level.
As shown in Figure 5, unlike the token-level Action MoE, this variant averages the entire token sequence before feeding it into the router, effectively performing expert selection along the batch (trajectory) dimension rather than the token dimension. Each trajectory is treated as a single entity representing a specific scenario or driving skill.
Moreover, to explicitly guide our model toward meaningful skill specialization—mirroring structured and intuitive human-defined skill categories—we utilize driving skill labels $y_{k}\in\{1,\dots,K\}$, annotated based on scenarios, and train the skill router via a cross-entropy loss as well:
@@ 2.4.2 Trajectory-level ActionMoE
Additionally, we optimize the entire Action MoE module using a flow-matching trajectory loss $\mathcal{L}_{\text{FM}}$ to ensure accurate trajectory predictions:
@@
where $\lambda_{1}$ represents loss weight of flow matching policy, $\lambda_{2}$ represents loss weight of action router. We introduce noise into the action router following [8], which increases randomness and encourages exploration, effectively mitigating the risk of expert collapse.
@@ 2.5 Two Stage Training: From Teacher-Forcing to Adaptive Training
DriveMoE loads the Paligemma VLM pretrained weights [2] and we finetune it in the autonomous driving scene via a two-stage training procedure. In the first stage, both vision and action MoEs only select ground-truth experts while the router is trained jointly, which significantly stabilize the training. In the second stage, we transition to select experts based on the outputs of Vision and Action MoE routers, removing reliance on GT annotation about experts. It develops robustness against potential mistakes or inaccuracies made by routers, thereby enhancing the overall model’s generalization capability under realistic inference conditions.
Table 2: Results on the Bench2Drive Benchmark(Closed-Loop and Open-Loop). * denotes expert feature distillation.
Method Venue Closed-loop Metric Open-loop Metric
[…]
TCP-traj* [54] NeurIPS 2022 59.90 30.00 76.54 18.08 1.70
AD-MLP [61] Arxiv 2023 18.05 0.00 48.45 22.63 3.64
VAD [31] ICCV 2023 42.35 15.00 157.94 46.01 0.91
UniAD-Base [16] CVPR 2023 45.81 16.36 129.21 43.58 0.73
ThinkTwice* [25] CVPR 2023 62.44 31.23 69.33 16.22 0.95
DriveAdapter* [22] ICCV 2023 64.22 33.08 70.22 16.01 1.01
GenAD [63] ECCV 2024 44.81 15.90 - - -
DriveTrans [28] ICLR 2025 63.46 35.01 100.64 20.78 0.62
MomAD [48] CVPR 2025 44.54 16.71 170.21 48.63 0.82
WoTE [36] ICCV 2025 61.71 31.36 - - -
DriveMamba-L [49] ICLR 2026 66.82 37.73 152.91 18.77 0.70
DiffAD [52] Arxiv 2025 67.92 38.64 - - 1.55
Raw2Drive [58] NeurIPS 2025 71.36 50.24 214.17 22.42 -
[…]
@@ 3.1 Datasets & Benchmark & Metric
We employ the CARLA simulator [10] (version 0.9.15.1) for closed-loop driving performance evaluation, and adopt the latest public closed-loop evaluation benchmark, Bench2Drive [27] which includes 220 routes with one challenging corner case per route for analysis of different driving abilities. It provides an official training set, where we use the base set (1000 clips, 950 training, 50 test/validation) for fair comparison with all the other baselines.
Table 3: Ablation study on Vision MoE. Compare different camera view combinations and supervision signals. $F$, FLFL, FRFR, and $B$ indicate the front, front-left, front-right, and back views, respectively, while BLBL and BRBR represent the back-left and back-right views. Fixed View means selecting a specific view. Dynamic View refers to the camera view dynamically selected by the vision router as the top-1 relevant view according to scene context. Exp 1 denotes our baseline Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$, which models surrounding agents’ velocities from two consecutive front-view images, and Exp 9 denotes DriveMoE, which adds a dynamically selected view with explicit supervision to enhance perception learning. Memory is evaluated at batch size=1. All experiments use the previous-frame front view by default.
@@
1 $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ Fixed - 55.85 30.00 100ms 4100
2 $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ Fixed - 62.38 33.64 260ms 5100
3 $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ Fixed - 61.52 32.73 260ms 5100
4 $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ Fixed - 63.26 31.82 260ms 5100
5 $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ Fixed - 64.92 33.64 400ms 7400
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ Fixed - 64.18 33.64 550ms 9600
7 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Fixed - 62.27 31.36 700ms 11800
8 $\checkmark$ - - - - - Dynamic $\times$ 69.71 44.09 260ms 5100
@@
3.2 Comparison with State-of-the-Art Methods
As shown in Table 2, our proposed method achieves state-of-the-art (SOTA) performance in terms of both driving score and success rate on the Bench2Drive closed-loop benchmark. Specifically, compared to the baseline Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$, our method improves the driving score by 22.8% and the success rate by 62.1%. On the open-loop metric, our method achieves the lowest L2 error. We observe that diffusion policy-based trajectory prediction significantly reduces L2 errors compared to traditional methods. However, as highlighted in prior studies such as AD-MLP [60], TransFuser++ [21], and Bench2Drive [27], open-loop metrics mainly serve as indicators of model convergence, whereas closed-loop metrics provide a more reliable assessment of true driving performance. Moreover, in the multi-dimensional capability evaluation, as shown in Table 1, our method obtains state-of-the-art results across five key capabilities and their overall average.
We use the official 220 routes and official metrics of Bench2Drive for evaluation. The Driving Score (DS) is defined as the product of Route Completion and Infraction Score, capturing both task completion and rule adherence. The Success Rate (SR) measures the percentage of routes completed successfully within the allocated time and without committing any traffic violations. Efficiency quantifies the vehicle’s velocity relative to surrounding traffic, encouraging progressiveness without aggression. Comfort reflects the smoothness of the driving trajectory. Meanwhile, Bench2Drive evaluates driving capabilities across multiple critical dimensions, including tasks such as Merging, Overtaking, Emergency Braking, Yielding, and Traffic Signs. All results are averaged over three runs.
@@ 3.3 Ablation Study
Vision MoE. As shown in Table 3, we investigate the contribution of camera view selection and supervision signals within our Vision MoE module. The baseline (Exp 1, Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$) utilizes two consecutive front-view images primarily to estimate velocities of surrounding agents. Adding a third fixed view such as the back view (Exp 4), front-left view (Exp 2), or front-right view (Exp 3) provides additional spatial context, yielding moderate improvements. According to Exp 2-4, adding individual additional views provides more complementary perceptual information, leading to improved performance. However, when multiple views are added simultaneously (as in Exp 5-7), the number of visual tokens increases significantly, which introduces greater training complexity and convergence difficulty. As a result, model performance degrades, and the latency increases.
@@
By introducing dynamically selected views without supervision (Exp 8), the driving score and success rate significantly improve. Ultimately, incorporating explicit supervision signals (Exp 9, DriveMoE) for the dynamic view selection further enhances both driving score and success rate, demonstrating the effectiveness of our Vision MoE module in leveraging dynamic and supervised multi-view perception. Table 4 shows the accuracy of the vision and action routers on the test set under open-loop evaluation.
Token vs Trajectory Level Action MoE. As shown in Table 5, we compare the token-level and trajectory-level routers, and observe that the trajectory-level router consistently outperforms the token-level counterpart. So We use Trajectory Level Action MoE as our default Action MoE.
Action MoE. For clarity, the term Action MoE in the following sections specifically refers to the Trajectory-Level Action MoE. We investigate the impact of the number of non-shared experts within our Action MoE, as shown in Table 6. Specifically, Exp 1 corresponds to the original five skills defined by Bench2Drive [27], while Exp 2 introduces an additional expert for the classic ParkingExits scenario, resulting in improved performance. To further analyze the effect of expert specialization, we conducted additional experiments: Exp 2–3 compare action supervision strategies, where the action supervision yields noticeable improvement; Exp 4-5 introduces additional shared experts, whose performance remains comparable to the Exp 2; Moreover, Exp 6 adds experts targeting several challenging scenarios identified from Exp 2, and Exp 7 assigns a distinct expert to each of the 44 scenarios in Bench2Drive. We observe that excessively increasing the number of experts (Exp 5-6) negatively affects performance due to the induced load imbalance among experts. Thus, an appropriate balance in the number of specialized experts is crucial for optimal driving performance.
@@
Table 5: Token vs Trajectory Level Action MoE.
[…]
Token-Level 3 6 65.62 32.27
Traj-Level 3 6 73.88 48.64
[…]
1 1 5 $\checkmark$ 73.81 47.73
[…]
3 1 6 $\times$ 70.38 45.00
4 3 6 $\checkmark$ 73.88 48.64
5 5 6 $\checkmark$ 73.46 47.73
6 1 13 $\checkmark$ 70.88 44.50
7 1 44 $\checkmark$ 68.22 43.18
Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ vs DriveMoE. We conduct ablation studies to evaluate the individual contributions of the Vision MoE and Action MoE within our DriveMoE framework. As shown in Table 7, removing either the Vision MoE or the Action MoE leads to a noticeable decline in both driving score and success rate. Compared to the $\boldsymbol{\pi}_{\boldsymbol{0}}, Drive$MoE substantially improves driving performance, highlighting the complementary effectiveness of both MoE modules.
Table 7: Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ vs DriveMoE. Evaluate the Vision MoE and Action MoE. ”w/o” denotes removing the respective modules.
[…]
Drive-$\boldsymbol{\pi}_{\boldsymbol{0}}$ 55.85 30.00
w/o Vision MoE 68.68 42.45
w/o Action MoE 67.31 40.56
DriveMoE 74.22 48.64